In a series LCR circuit with L = 5/3 H, R = 10 Ω and C = 1/30 F and a source of emf
E = 600 V, determine the charge on the capacitor as a function of time. It is given that
the initial charge on the capacitor is zero and the initial current in the circuit is 9A.
1
Expert's answer
2019-09-09T11:10:48-0400

The RLC circuit is descrided by the initial value problem


"L\frac{d^2 q}{dt^2}+R\frac{d q}{dt}+\frac{1}{C}q=E"

"q(0)=q_0,\quad I(0)=q'(0)=I_0"


where "q(t)" is a charge on the capacitor.

In our case we have


"\frac{5}{3}\frac{d^2 q}{dt^2}+10\frac{d q}{dt}+30q=600"

or

"\frac{d^2 q}{dt^2}+6\frac{d q}{dt}+18q=360"

and

"q(0)=0,\quad I(0)=q'(0)=9"

Solution of the IVP

"q(t)=e^{-3t}(20e^{3t}-20\cos(3t)-17\sin(3t))"